letisia:

Write the equation of the line, in standard form, that has a y-intercept of 2 and is parallel to 2x + y = -5. Include your work in your final answer. Type your answer in the box provided submit your solution.

Mercury:

2x + y = -5 subtracting 2x from both sides y = -2x - 5 comparing this to y = mx + b, the slope (m-value) is -2 since we want an equation that is *parallel* to this line, our new equation also should have slope -2 the desired y-intercept (a.k.a b-value) is 2 plug in the appropriate m and b values into a new y = mx + b equation